Grade 12th passMechanics

To a 25ml H2O2 solution excess of acidified solution of KI was added. The iodine liberated required 20 ml of 0.3 N Na2S2O3 solution. The volume strength of H2O2 solution is

Sana khan
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
Meq of H2O2 = Meq of Na2S2O3 If N is normality of H2O2, then N × 25 = 0.3 × 20 ⇒ N = 0.24 ⇒ Volume strength = N × 5.6 = 1.334 V